OpenStudy (anonymous):

Please help me thank you to everyone.Given the two functions below, what is the x-intercept of g(x)? f(x) = x^2 g(x) = f(x –7)

OpenStudy (bibby):

the x intercept is a point on the graph where y=0,it's of the format (x,0) start by finding out what f(x-7) is

OpenStudy (anonymous):

thanks bibby

OpenStudy (bibby):

so, what's t he answer?

OpenStudy (anonymous):

7?

OpenStudy (bibby):

\(g(x) = f(x –7)=(x-7)^2=(x-7)(x-7)=x^2-14x+49\\g(x)=x^2-14x+49\) then we set y equal to 0

OpenStudy (bibby):

god dammit

OpenStudy (bibby):

\(g(x) = f(x-7)=(x-7)^2=(x-7)(x-7)=x^2-14x+49\\g(x)=x^2-14x+49\)

OpenStudy (bibby):

so x^2-14x=0 I'm not sure why I did this so redundantly but yes x=7

OpenStudy (anonymous):

okay so what do i do now @bibby

OpenStudy (bibby):

you're done, the x intercept is at x=7 or (7,0)

OpenStudy (anonymous):

thank you so much!!

OpenStudy (bibby):

anytime